Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: C config; intense observing pressure in the 1200-2100 LST range - proposed to default to science observing on non-maintenance days to begin ~1500 LST to help alleviate the pressure (over the next few weeks); double maintenance day (Tue-Wed); Thanksgiving operations shutdown/re-start went smoothly. Call for proposals draft this week. PSC/RSRO committee meet next week.
  • Key dates:
  • Staff Travel:
    • Vivek away until 02 Dec
    • Rick, Michael away until 03 Dec
  • Visitors:
    • Bob Sault

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon: Testing
  • Tue, Wed: Maintenance days
  • Thu-Fri: 0800-1300 MST for testing/development; science 1300+
  • Interrupt status for additional 3-bit testing; expected to be lab work only this week

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

Commissioning Reports

Imaging: McMullin (for Liu/Ott): Proto-Mosaicing setup tool

RFI: Venkata/McMullin: L-band bandpass oscillation.

Group updates: *; User Support, Scheduling, High Frequency, Algorithm

Commissioning Group updates

In preparation

  • * Flagging*: Venkata: Autoflagging in CASA (Coming soon)
  • Polarization: Myers: Multi-band leakage/PA
  • Array: Dhawan/Rupen: RF switch tests with rapid monitoring; issue still affecting; being replaced as identified. Looking at acquiring new switches to solve the problem in the long term.

Commissioning Test Planning

  • Antenna: Ongoing Ka, Q band wobble investigation (McMullin/Long)
  • Correlator: CM 2010-11-29 20:33 UT version checkout (McMullin/Sowinski)
  • Antenna: EA26 checkout for array use (Sowinski)
  • M&C: MCAF version update testing (McMullin)
  • High Frequency: Claussen: Weak calibrator tests; needed.
  • Spectral Line: Momjian: Zeeman followup
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Postponed until wideband X available on EA24.

RSRO/ECSO Reports

Misc:

AOB

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 2010-11-28
Topic revision: r4 - 2010-11-30, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback